# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v2

@@ 6 Conclusion
In this paper, we have proposed two algorithms for joint prior selection and regret minimization in GP bandits based on GP-TS. We have analyzed the algorithms theoretically and have experimentally evaluated both algorithms on synthetic and real-world data. We find that they both select the true prior more often and obtain lower regret than previous work due to lowering the amount of optimistic exploration.
@@ A.1 PE-GP-TS
Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$. Additionally, for Eq.˜5 we must take an additional union bound over $p\in P$.
@@
$\mathbb{P}\left(\left|\frac{\tilde{f}_{t,p}(x)-\mu_{t,p^{*}}(x)}{\sigma_{t,p^{*}}(x)}\right|>\sqrt{\beta_{t}}\right)$ $\leq\exp(-\beta_{t}/2)$ (12)
@@
Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Eq.˜5 holds w.p. at least $1-\delta/2$. By the same reasoning and skipping the union bound over $P$, Eq.˜4 holds w.p. at least $1-\delta/2$. Thus, both events hold w.p. at least $1-\delta$. ∎
@@
Using Eqs.˜14 and 21, we can bound the cumulative regret as follows:
@@
Finally, applying Eq.˜15, we obtain the result
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (33)
@@ A.2 UCB-analysis of HP-GP-TS
Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]=\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$. In our case, note that $f(x)|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}}(x),\sigma^{2}_{t,p^{*}}(x))$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$. Hence,
@@ A.3 Analysis of expected posterior probabilities
Note that $N_{p}({\mathbf{y}})/N_{\tilde{p}}({\mathbf{y}})=\exp\left(-\frac{1}{2}\left({\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}-({\mathbf{y}}^{T}-{\bm{\mu}}){\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})\right)\right)=:\exp(Q({\mathbf{y}}))$ and that the quotient in the integrand can be written as $q(x)=\frac{e^{x}c}{e^{x}c+(1-c)}$ for $c:=P_{0}(p)\in(0,1)$. Since we wish to lower bound the expectation of $P_{t+1}(p)$, we will utilize the lower bounds $q(x)\geq ce^{x}$ $\forall x\leq 0$ and $q(x)\geq 1-\frac{1-c}{c}e^{-x}$ $\forall x\in{\mathbb{R}}$. The first lower bound will be applied to the region $H:=\{{\mathbf{y}}\big|Q({\mathbf{y}})\leq 0\}$ whilst the second is applied to the complement $H^{c}$:
@@
Note that a sharp bound of Eq.˜77 in terms of elementary functions can be obtained using bounds on the error function (Cook, 2018; Abramowitz & Stegun, 1972):
@@
A.4 Information-theoretic regret bound for GP-TS
We begin by showing that the rewards are subgaussian. {lemma}[] Fix $x\in{\mathcal{X}}$ and let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$. If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ with $k_{1,p^{*}}(\cdot,\cdot)\leq\sigma_{0}^{2}$, then $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian.
Proof.
Start by considering $Z=\sum_{p\in P}\mathds{1}\{p^{*}=p\}(f_{p}(x)-\mu_{t,p}(x))$ where $f_{p}\sim{\mathcal{GP}}_{p}$ such that $f_{p}(x)-\mu_{t,p}(x)$ is $\sigma_{0}$-subgaussian. Then,
${\mathbb{E}}[\lambda Z|H_{t}]$ $={\mathbb{E}}[{\mathbb{E}}[\exp(\lambda Z)\,|\,p^{*}=p,H_{t}]$ (80)
$=\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p){\mathbb{E}}[\exp(\lambda(f_{p}(x)-\mu_{t,p}(x)))\,|\,p^{*}=p,H_{t}]$ (81)
$\leq\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p)\exp(\sigma_{t,p}^{2}(x)\lambda^{2}/2)$ (82)
$\leq\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p)\exp(\sigma_{0}^{2}\lambda^{2}/2)$ (83)
$=\exp(\sigma^{2}_{0}\lambda/2)$ (84)
Hence, $Z$ is $\sigma$-subgaussian. Then, $Z+\epsilon_{t}|H_{t}$ is $\sqrt{\sigma^{2}_{0}+\sigma^{2}}$-subgaussian since $Z|H_{t}$ and $\epsilon_{t}|H_{t}$ are independent. ∎
Then, we state and prove the information-theoretic regret bound for HP-GP-TS. See 4.3
Proof.
Here, we analyze the regret of HP-GP-TS using the information-theoretic framework of Russo & Van Roy (2016). The proof presented here follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise. The idea of the information-theoretic framework is to express the regret as follows
${\mathbb{E}}\left[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\right]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}{\mathbb{E}}\big[\underbrace{f(x^{*})-f(x_{t})}_{\geq 0}|H_{t}\big]\right]$ $(\text{Tower rule})$ (85)
$={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}\left[f(x^{*})-f(x_{t})|H_{t}\right]^{2}\cdot\frac{I(\cdot\,;\cdot\,|H_{t})}{I(\cdot\,;\cdot\,|H_{t})}}\right]$ (86)
where $I(\cdot\,;\cdot\,|H_{t})$ is the mutual information between two carefully chosen variables such that ${\mathbb{E}}\left[f(x^{*})-f(x_{t})|H_{t}\right]^{2}\leq CI(\cdot\,;\cdot\,|H_{t})$ for some $C$.
Let ${\mathbb{E}}_{t}[\cdot]={\mathbb{E}}[\cdot|H_{t}],{\mathbb{P}}_{t}[\cdot]={\mathbb{P}}[\cdot|H_{t}]$. Then,
${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{E}}_{t}[f(x)|x^{*}=x]-\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x){\mathbb{E}}_{t}[f(x)|x_{t}=x]$ (87)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)|x_{t}=x]\right)\quad$ $\left({\mathbb{P}}_{t}(x_{t}=x)={\mathbb{P}}_{t}(x^{*}=x)\right)$ (88)
@@ A.4 Information-theoretic regret bound for GP-TS
Let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$. Note that $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian. Consequently, $\log{\mathbb{E}}_{t}[\exp(\lambda Z)]\leq\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ $\forall\lambda\in{\mathbb{R}}$.
Fix $x^{*},x\in{\mathcal{X}}$. Then,
$\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)-\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ (90)
$\leq\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)$ (91)
$\quad-\log{\mathbb{E}}_{t}[\exp(\lambda(f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)])]$ $(Z|H_{t}\text{ is }\sqrt{\sigma_{0}^{2}+\sigma^{2}}\text{-subgaussian})$ (92)
$=\lambda{\mathbb{E}}_{t}[Z|x^{*}=x^{*}]-\log{\mathbb{E}}_{t}[\exp(\lambda Z)]$ (93)
[…]
Now, let $\lambda=\frac{1}{\sigma_{0}^{2}+\sigma^{2}}\left({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)$, then the following holds for all $x^{*},x\in{\mathcal{X}}$
${\mathbb{E}}_{t}\left[f(x)|x^{*}=x^{*}\right]-{\mathbb{E}}_{t}[f(x)]\leq\sqrt{2(\sigma_{0}^{2}+\sigma^{2})D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)}.$ (95)
Let $I_{t}(\cdot\,;\cdot\,)=I(\cdot\,;\cdot\,|H_{t}=h_{t})$. Then,
$I_{t}(x^{*};(x_{t},y_{t}))=I_{t}(x^{*};x_{t})+I_{t}(x^{*};y_{t}|x_{t})$ $(\text{Chain rule})$ (96)
$=I_{t}(x^{*};y_{t}|x_{t})$ $(x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (97)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};y_{t}|x_{t}=x)$ (98)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x_{t})+\epsilon_{t}|x_{t}=x)$ (99)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x)+\epsilon_{t})$ $(f,x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (100)
$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)\Bigg(\sum_{x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (101)
$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Bigg)$ (102)
$=\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ $(x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t})$ (103)
$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)$ (104)
Putting the above together, we now bound ${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}$:
${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}=\left(\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)\right)^{2}$ (105)
$\leq|{\mathcal{X}}|\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)^{2}\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ $(\text{Cauchy-Schwarz})$ (106)
$\leq|{\mathcal{X}}|\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\left({\mathbb{E}}_{t}[f(x)|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ (107)
$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})\sum_{x,x^{*}\in{\mathcal{X}}}\Big({\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (108)
$\qquad\qquad\qquad\qquad\qquad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Big)$ (109)
$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})I_{t}(x^{*};(x_{t},y_{t})).$ (110)
Returning to the full regret,
${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{t}\left[f(x^{*})-f(x_{t})\right]^{2}\cdot\frac{I_{t}(x^{*}\,;(x_{t},y_{t}))}{I_{t}(x^{*}\,;(x_{t},y_{t}))}}\right]$ (111)
$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}{\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (112)
Then, the expectation can then be bounded as
${\mathbb{E}}\bigg[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\bigg]=$ $\leq{\sum_{t\in[T]}}{\mathbb{E}}\left[\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (114)
$\leq{\sum_{t\in[T]}}1\cdot\sqrt{{\mathbb{E}}\left[I_{t}(x^{*}\,;(x_{t},y_{t}))\right]}$ $\left(\text{Jensen's inequality}\right)$ (115)
$={\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{H_{t}}\left[I(x^{*}\,;(x_{t},y_{t})|H_{t}=h_{t})\right]}$ (116)
$={\sum_{t\in[T]}}\sqrt{I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ (117)
$\leq\sqrt{{\sum_{t\in[T]}}1^{2}{\sum_{t\in[T]}}I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ $(\text{Cauchy-Schwarz})$ (118)
Finally, the regret can be bounded as
${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $=\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TI(x^{*};H_{T})}$ (119)
$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TH(x^{*})}$ (120)
$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{T\log|{\mathcal{X}}|}$ (121)
Appendix B Description of kernels
[…]
Appendix C Additional experimental details

### v3

@@ 6 Conclusion
In this paper, we have proposed two algorithms for adaptive prior selection and regret minimization in GP bandits based on GP-TS. We have analyzed the algorithms theoretically and have experimentally evaluated both algorithms on synthetic and real-world data. We find that they both select the true prior more often and obtain lower or comparable regret than previous work due to lowering the amount of optimistic exploration.
@@ A.1 PE-GP-TS
Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$. Additionally, for Equation˜5 we must take an additional union bound over $p\in P$.
@@
$\mathbb{P}\left(\left|\frac{\tilde{f}_{t,p}(x)-\mu_{t,p}(x)}{\sigma_{t,p}(x)}\right|>\sqrt{\beta_{t}}\right)$ $\leq\exp(-\beta_{t}/2)$ (13)
@@
Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Equation˜5 holds w.p. at least $1-\delta/2$. By the same reasoning and skipping the union bound over $P$, Equation˜4 holds w.p. at least $1-\delta/2$. Thus, both events hold w.p. at least $1-\delta$. ∎
@@
Using Equations˜15 and 22, we can bound the cumulative regret as follows:
@@
Finally, applying Equation˜16, we obtain the result
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{T\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (34)
@@ A.2 UCB-analysis of HP-GP-TS
Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]\leq\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$. In our case, note that $f(x)|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}}(x),\sigma^{2}_{t,p^{*}}(x))$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$. Hence,
@@ A.3 Analysis of expected posterior probabilities
Note that $N_{p}({\mathbf{y}})/N_{\tilde{p}}({\mathbf{y}})=\exp\left(-\frac{1}{2}\left({\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}-({\mathbf{y}}-{\bm{\mu}})^{T}{\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})\right)\right)=:\exp(Q({\mathbf{y}}))$ and that the quotient in the integrand can be written as $q(x)=\frac{e^{x}c}{e^{x}c+(1-c)}$ for $c:=P_{0}(p)\in(0,1)$. Since we wish to lower bound the expectation of $P_{t+1}(p)$, we will utilize the lower bounds $q(x)\geq ce^{x}$ $\forall x\leq 0$ and $q(x)\geq 1-\frac{1-c}{c}e^{-x}$ $\forall x\in{\mathbb{R}}$. The first lower bound will be applied to the region $H:=\{{\mathbf{y}}\big|Q({\mathbf{y}})\leq 0\}$ whilst the second is applied to the complement $H^{c}$:
@@
Note that a sharp bound of Equation˜78 in terms of elementary functions can be obtained using bounds on the error function (Cook, 2018; Abramowitz and Stegun, 1972):
@@
Appendix B Technical issues with MixTS regret bound in the linear setting
Theorem 1 in Hong et al. (2022a) provides a regret bound for MixTS in the linear setting. The linear setting assumes that the true parameter $\theta^{*}\sim{\mathcal{N}}(\theta_{0,S_{*}},\Sigma_{0,S_{*}})$ where the latent state $S_{*}$ is sampled from a discrete mixture prior $P_{0}$. The proof of Theorem 1 in Hong et al. (2022a) contains a few non-obvious steps that seem difficult to motivate. We use the notation of Hong et al. (2022a).
  1. 1.
    Eq. (9) in Hong et al. (2022a) uses the TS property that the true prior and optimal arm is equal in distribution to the selected prior and selected arm given the history: $A_{t,*}^{\top}\bar{\theta}_{t,S_{*}}|H_{t}\overset{d}{=}A_{t}^{\top}\bar{\theta}_{t,S_{t}}|H_{t}$ (equivalent to $\mu_{t,p^{*}}(x^{*})|H_{t}\overset{d}{=}\mu_{t,p_{t}}(x_{t})|H_{t}$ in our notation). However, Eq. (9) additionally conditions on the event $E_{0}=\{\|\theta_{*}-\theta_{0,S_{*}}\|_{\Sigma_{0,S_{*}}^{-1}}\leq\sqrt{2d\log(dn)}\}$ where $\theta_{*}$ lies close to its prior mean. The TS property does not hold under this event since it modifies the distribution of the linear parameter $\theta^{*}$ but not the sampled parameters $\theta_{t}$, thus changing the distribution of the optimal arm $A_{t,*}$ but not the selected arm $A_{t}$. Consider the example in Figure˜8, if $E_{0}$ holds then $\theta_{*}$ lies in the blue regions and thus $a_{2}$ is optimal w.p. 0. If $E_{0}^{c}$ holds, then $a_{2}$ is optimal with a non-zero probability. However, MixTS is oblivious to $E_{0}$ given the history and thus $A_{t,*}|H_{t},E_{0}\not\overset{d}{=}A_{t}|H_{t}$. This counterexample illustrates the overall idea but we have not validated that the scale of the arms and the blue regions are feasible.
  2. 2.
    Five lines above Eq. (9) in Hong et al. (2022a), it is stated that the regret is upper-bounded by a constant $M$ whenever $E_{0}$ occurs. However, from Eq (9) to the first term in step 3 of their analysis (page 15), the bound of $M$ is applied implicitly to $A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*}$ without motivation. For the setting with bounded rewards, then $\bar{\theta}_{t,S_{t}}$ is also bounded but for Gaussian rewards $\bar{\theta}_{t,S_{t}}$ can be unbounded.
  3. 3.
    The second term in Eq. (9) contains the indicator function $\mathbf{1}\{E_{0}\}$: $\mathbb{E}[(A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*})\mathbf{1}\{E_{0}\}]$. In step 3 (page 15), this indicator function is dropped without motivation: $\mathbb{E}[\langle A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*}\rangle_{M}]$ where $\langle\cdot\rangle_{M}=\min(\cdot,M)$ for the bound $M$. If the expression inside is non-negative w.p. 1, then this step would be valid but this is not the case.
  4. 4.
    From our understanding, the final equation on page 15 adds and subtracts the confidence bound and adds a zero-mean Gaussian inside a minimum. However, adding a zero-mean Gaussian inside a minimum reduces the expectation but the analysis seems to assume that it would increase the expectation. I.e. it is seemingly assumed that $\mathbb{E}[\min(M,X)]\leq\mathbb{E}[\min(M,X+\epsilon_{t})$ for a constant M and random variable $X$. However, the reverse inequality is true.
Theorem 1 of Hong et al. (2022a) provides a novel proof technique that could potentially be applied to HP-GP-TS. But the technical issues presented above present a few challenges.
Figure 8: Potential counterexample to Eq (9) in Hong et al. (2022a). The blue regions represent when the event $E_{0}$ holds and the black dots represent the two arms $a_{1}$ and $a_{2}$.
Appendix C Extending the analysis of MixTS to HP-GP-TS
In this section, we explore how the analysis of Hong et al. (2022a) can be applied to the Gaussian process setting considered in this paper. The core idea of the analysis of Hong et al. (2022a) is to introduce confidence sets for the priors (or latent variables in their terminology). The confidence intervals closely resemble the prior elimination criteria introduced by Ziomek et al. (2025). In our notation, the equivalent confidence set $C_{t}$ of Hong et al. (2022a) is defined as
$C_{t}=\left\{p\in P:\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(x_{s})-\sqrt{\beta_{T}}\sigma_{s,p_{s}}(x_{s})-y_{s})\leq 2\sigma\sqrt{N_{t}(p)\log T}\right\}.$ (81)
We define $G_{t}(p)=\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(x_{s})-\sqrt{\beta_{T}}\sigma_{s,p_{s}}(x_{s})-y_{s})$. The prior elimination criteria for PE-GP-UCB and PE-GP-TS can be expressed as a confidence set in the following way:
$\tilde{C}_{t}=\Bigg\{p\in P:\forall s<t\quad\Bigg|\sum_{\begin{subarray}{c}i\leq s\\ p_{i}=p\end{subarray}}y_{i}-\mu_{i,p}(x_{i})\Bigg|\leq\sqrt{2\sigma^{2}N_{t}(p)\log(t^{2}|P|\pi^{2}/3)}+\sum_{\begin{subarray}{c}i\leq s\\ p_{i}=p\end{subarray}}\sqrt{\beta_{i}}\sigma_{i,p}(x_{i})\Bigg\}.$ (82)
The main difference between $C_{t}$ and $\tilde{C}_{t}$ is that $\tilde{C}_{t}$ is monotonically decreasing in cardinality since eliminated priors cannot be selected by PE-GP-UCB and -TS. Additionally, $\tilde{C}_{t}$ uses a symmetric criteria and the confidence parameters in $C_{t}$ use the final time step $T$ rather than the current time step $t$.
To apply the analysis of Hong et al. (2022a) directly to a GP-setting would require resolving the issues discussed in Appendix˜B. To circumvent some of these issues, we attempt to analyze the regret of HP-GP-TS by directly applying a UCB-decomposition and manipulate the decomposition such an expression similar to $G_{t}(p_{t})$ appears.
Recall, the regular UCB-decomposition to the Bayesian cumulative regret:
${\text{BR}}(T)$ $={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-f(x_{t})]$ (83)
$={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})+U_{t,p^{*}}(x^{*})-f(x_{t})]$ (84)
$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p_{t}}(x_{t})-f(x_{t})]}_{(2)}\quad(p_{t},x_{t}|H_{t}\overset{d}{=}p^{*},x^{*}|H_{t})$ (85)
By the same steps as in the proof of Theorem˜4.2, $(1)\leq\pi^{2}/6$. Hence, we focus on the second term and condition on whether the selected prior lies in the confidence set:
$(2)={\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\in C_{t}\}]+{\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\notin C_{t}\}].$ (86)
To bound the right-term above, we would like $(U_{t,p_{t}}(x_{t})-f(x_{t}))$ to be bounded such that it is sufficient to analyze the probability $\mathds{1}\{p_{t}\notin C_{t}\}$. However, we have not identified a method of doing so. We continue to study the left term. Since the noise $\epsilon_{t}$ is independent of $\mathds{1}\{p_{t}\in C_{t}\}$ and has zero mean, we can add it inside the expectation and add-and-subtract the confidence radius:
${\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\in C_{t}\}]$ (87)
$={\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (88)
$={\sum_{t\in[T]}}{\mathbb{E}}[(\mu_{t,p_{t}}(x_{t})-\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (89)
$\quad+{\sum_{t\in[T]}}2{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (90)
$\leq\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[(\mu_{t,p_{t}}(x_{t})-\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]}_{(3)}$ (91)
$\quad+{\sum_{t\in[T]}}2{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})]$ $(\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\geq 0)$ (92)
Note that the sum of $\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})$ can be bounded in terms of the worst-case MIG $\hat{\gamma}_{T}$} by Equation˜16. To bound (3), following the argument of Hong et al. (2022a), we note that for each prior $p\in P$ there must be some last time $t^{\prime}(p)\leq T$ that $p$ was in the confidence set and chosen by the algorithm: $p\in C_{t^{\prime}}$ and $p=p_{t^{\prime}}$. Since $p\in C_{t^{\prime}}$, then $G_{t^{\prime}}(p)\leq 2\sigma\sqrt{N_{t^{\prime}}(p)\log T}$. Therefore, for each prior $p$, we can split the sum into $G_{t^{\prime}}(p)$ and the final term at time $t^{\prime}$:
$(3)$ $\leq{\mathbb{E}}\bigg[\sum_{p\in P}(G_{t^{\prime}}(p)+(\mu_{t^{\prime},p}(x_{t^{\prime}})-\sqrt{\beta_{t^{\prime}}}\sigma_{t^{\prime},p}(x_{t^{\prime}})-f(x_{t^{\prime}})-\epsilon_{t^{\prime}}))\mathds{1}\{p\in C_{t^{\prime}}\}\bigg]$ (93)
$\leq{\mathbb{E}}\bigg[\sum_{p\in P}2\sigma\sqrt{N_{t^{\prime}}(p)\log T}\bigg]$ (94)
$\quad+{\mathbb{E}}\bigg[\sum_{p\in P}(\mu_{t^{\prime},p}(x_{t^{\prime}})-\sqrt{\beta_{t^{\prime}}}\sigma_{t^{\prime},p}(x_{t^{\prime}})-f(x_{t^{\prime}})-\epsilon_{t^{\prime}}))\mathds{1}\{p\in C_{t^{\prime}}\}\bigg].$ (95)
The first sum above can be bounded using that $\sum_{p\in P}N_{T}(p)=T$ and Jensen’s inequality. The major challenge is to show that the final term is bounded. This requires creating an event s.t. the terms inside are bounded with high probability. However, as the issues in Appendix˜B highlight, if the event is in introduced at an earlier step then it must not invalidate the subsequent manipulations.
Appendix D Description of kernels
@@
Appendix E Additional experimental details